OpenStudy (anonymous):

How do you write polar number. Say V is a polar number. Do you bold it? I thought bold is for vector?

OpenStudy (anonymous):

\( \bf{V} \) ?

OpenStudy (anonymous):

V=

OpenStudy (anonymous):

That vector isn't it? I am asking how the textbook represent a polar number. polar number is a number with angle.

OpenStudy (anonymous):

Do you write tilde below V?

OpenStudy (anonymous):

vector is bold and has arrow above.

OpenStudy (anonymous):



OpenStudy (unklerhaukus):

when you say a polar number , do you mean a vector represented in polar form?

OpenStudy (hba):

This question is 4 months old :o

OpenStudy (unklerhaukus):

\[\vec z={\bf z} =(a,b)=a{\bf\hat x}+b{\bf\hat y}=a{\textbf e}_1+b{\textbf e}_2\] \[\vec z={\bf z} =(r,\theta)=re^{i\theta}=r\big(\cos(\theta)+i\sin(\theta)\big)=r\hat{\textbf r}\]